# Two-dimensional delta Bose gas in a weighted space

### v1

@@ Acknowledgment
$\int_{0}^{\infty}\mathrm{d}s\,e^{-s}/\sqrt{s^{2}+ss_{0}}=-2\log\sqrt{s_{0}}+2\int_{0}^{\infty}\mathrm{d}s\,e^{-s}\log(\sqrt{s}+\sqrt{s+s_{0}})$, and use (1.2) to simplify the result. Doing so gives
$C_{\varepsilon}(\lambda)=1-\frac{\beta_{\varepsilon}}{4\pi}(\log\lambda-\gamma_{\varepsilon}-D(c_{0}\varepsilon^{2}\lambda))\ ,$ (3.18)
where $\gamma_{\varepsilon}$ is a real constant such that $\gamma_{\varepsilon}\to\gamma\in\mathbb{R}$ as $\varepsilon\to 0$, and
$D(\lambda):=2\int_{0}^{\infty}\mathrm{d}s\,e^{-s}\log(\sqrt{s}+\sqrt{s+\lambda})\ .$ (3.19)
Put $c_{1}:=e^{\gamma+D_{\varepsilon}(0)+1}$. By (3.18), for all real $\lambda>c_{1}$ and small enough $\varepsilon$, $C_{\varepsilon}(\lambda)<1$. By (3.17), for all real $\lambda>0$, $C_{\varepsilon}(\lambda)>0$. Since the integral in (3.15) is a sum of convolutions, the Laplace transform of $B_{\varepsilon}(t)$ is a geometric series $\int_{0}^{\infty}\mathrm{d}t\,e^{-t\lambda}B_{\varepsilon}(t)=\beta_{\varepsilon}\sum_{k=1}^{\infty}C_{\varepsilon}(\lambda)^{k}=\beta_{\varepsilon}C_{\varepsilon}(\lambda)/(1-C_{\varepsilon}(\lambda)),$ for all $\lambda>c_{1}$. The last expression analytically continues to $z\in\mathbb{C}\setminus(-\infty,c_{1}]$. Given the Laplace transform, Mellin’s inversion formula yields $B_{\varepsilon}(t)=\int_{\supset}\frac{\mathrm{d}z}{2\pi\mathbf{i}}\,e^{zt}\beta_{\varepsilon}C_{\varepsilon}(z)/(1-C_{\varepsilon}(z))$, where :=$\supset\,:=\{(x-\mathbf{i}0)|-\infty<x\leq c_{1}\}\cup\{(x+\mathbf{i}0)|-\infty<x\leq c_{1}\}$ denotes the contour that goes from $\infty$- ∞ to $c_{1}$ along the lower side of the real axis and then from $c_{1}$ to $\infty$- ∞ along the upper side of the real axis. Add $\int_{\supset}\frac{\mathrm{d}z}{2\pi\mathbf{i}}\,e^{zt}\beta_{\varepsilon}$, which is zero, to the right-hand side. Doing so gives the contour-integral formula
@@
$F_{\varepsilon}(x)$ $:=\log|x|-\gamma_{\varepsilon}-\operatorname{Re}(D(-c_{0}\varepsilon^{2}x+\mathbf{i}0))\ ,$ (3.21)
$G_{\varepsilon}(x)$ $:=\big{(}\pi-\operatorname{Im}(D(-c_{0}\varepsilon^{2}x+\mathbf{i}0))\big{)}\mathbf{1}_{x>0}\ .$ (3.22)
Let $\log_{+}u:=\log(u\vee 1)$ and $u_{+}:=u\vee 0$ for $u\in\mathbb{R}$. It is not hard to check from (3.19) that, for all $x\in[-c_{1},\infty)$, $|G_{\varepsilon}(x)|\leq c$ and $F_{\varepsilon}(x)\geq\log|x|-c-\log_{+}(c_{0}\varepsilon^{2}x)$. Hence
$B_{\varepsilon}(t)\leq c\int_{-c_{1}}^{\infty}\mathrm{d}x\,\frac{e^{-tx}}{(\log(|x|/c_{2})-\log_{+}(c_{0}\varepsilon^{2}x))_{+}{}^{2}}\ ,$ (3.23)
@@
$\int_{-c_{1}}^{c_{2}e^{3}}+\int_{c_{2}e^{3}}^{(c_{2}e^{3})\vee(1/t)}+\int_{(c_{2}e^{3})\vee(1/t)}^{1/c_{0}\varepsilon^{2}}+\int_{1/c_{0}\varepsilon^{2}}^{\infty}$ (3.24)
@@
$I^{\prime}_{2}=\frac{(x-c_{2}e^{3})}{(\log x/c_{2})^{2}}\Big{|}^{(c_{2}e^{3})\vee(1/t)}_{c_{2}e^{3}}+\int_{c_{2}e^{3}}^{(c_{2}e^{3})\vee(1/t)}\mathrm{d}x\,\frac{2\,(x-c_{2}e^{3})}{x(\log x/c_{2})^{3}}\ .$ (3.25)
On the right-hand side, bound the terms by $t^{-1}(\log((e^{3})\vee(1/c_{2}t)))^{-2}$ and $\frac{2}{\log e^{3}}I^{\prime}_{2}$, respectively. Doing so gives $I^{\prime}_{2}\leq c\,e^{t}\,t^{-1}|\log(t\wedge\frac{1}{2})|^{-2}$. As for $I_{3}$, since $x\leq 1/\varepsilon^{2}c_{0}$, the denominator of the integrand is $|\log(x/c_{2})|^{2}$. This is $\geq|\log(t\wedge\frac{1}{2})|^{2}/c$ because $x\geq(c_{2}e^{3})\vee(1/t)$. For $I_{4}$, the denominator is $(\log\varepsilon^{2}-c)^{2}$, which is $\geq|\log(t\wedge\frac{1}{2})|^{2}/c$ because $t\leq\varepsilon^{2}c_{0}$. Bounding the denominators this way and releasing the range of integration to $x\in[0,\infty)$ give $I_{3}+I_{4}\leq c\,t^{-1}|\log(t\wedge\frac{1}{2})|^{-2}$.

### v2

@@ 3.1 Proving (30)
$\int_{0}^{\infty}\mathrm{d}s\,e^{-s}/\sqrt{s^{2}+ss_{0}}=-2\log\sqrt{s_{0}}+2\int_{0}^{\infty}\mathrm{d}s\,e^{-s}\log(\sqrt{s}+\sqrt{s+s_{0}})$, and use (2) to simplify the result. Doing so gives
$C_{\varepsilon}(\lambda)=1-\frac{\beta_{\varepsilon}}{4\pi}(\log\lambda-\eta_{\varepsilon}-D(c_{0}\varepsilon^{2}\lambda))\ ,$ (38)
where $\eta_{\varepsilon}$ is a real constant such that $\eta_{\varepsilon}\to\eta\in\mathbb{R}$ as $\varepsilon\to 0$, and
$D(\lambda):=2\int_{0}^{\infty}\mathrm{d}s\,e^{-s}\log(\sqrt{s}+\sqrt{s+\lambda})\ .$ (39)
Put $c_{1}:=e^{\gamma+D(0)+1}$. By (38), for all real $\lambda>c_{1}$ and small enough $\varepsilon$, $C_{\varepsilon}(\lambda)<1$. By (37), for all real $\lambda>0$, $C_{\varepsilon}(\lambda)>0$. Since the integral in (35) is a sum of convolutions, the Laplace transform of $B_{\varepsilon}(t)$ is a geometric series $\int_{0}^{\infty}\mathrm{d}t\,e^{-t\lambda}B_{\varepsilon}(t)=\beta_{\varepsilon}\sum_{k=1}^{\infty}C_{\varepsilon}(\lambda)^{k}=\beta_{\varepsilon}C_{\varepsilon}(\lambda)/(1-C_{\varepsilon}(\lambda)),$ for all $\lambda>c_{1}$. The last expression analytically continues to $z\in\mathbb{C}\setminus(-\infty,c_{1}]$. Given the Laplace transform, Mellin’s inversion formula yields $B_{\varepsilon}(t)=\int_{\supset}\frac{\mathrm{d}z}{2\pi\mathbf{i}}\,e^{zt}\beta_{\varepsilon}C_{\varepsilon}(z)/(1-C_{\varepsilon}(z))$, where :=$\supset\,:=\{(x-\mathbf{i}0)|-\infty<x\leq c_{1}\}\cup\{(x+\mathbf{i}0)|-\infty<x\leq c_{1}\}$ denotes the contour that goes from $\infty$- ∞ to $c_{1}$ along the lower side of the real axis and then from $c_{1}$ to $\infty$- ∞ along the upper side of the real axis. Add $\int_{\supset}\frac{\mathrm{d}z}{2\pi\mathbf{i}}\,e^{zt}\beta_{\varepsilon}$, which is zero, to the right-hand side. Doing so gives the contour-integral formula
@@
$F_{\varepsilon}(x)$ $:=\log|x|-\eta_{\varepsilon}-\operatorname{Re}(D(-c_{0}\varepsilon^{2}x+\mathbf{i}0))\ ,$ (41)
$G_{\varepsilon}(x)$ $:=\big{(}\pi-\operatorname{Im}(D(-c_{0}\varepsilon^{2}x+\mathbf{i}0))\big{)}\mathbf{1}_{x>0}\ .$ (42)
Let $\log_{+}u:=\log(u\vee 1)$ and $u_{+}:=u\vee 0$ for $u\in\mathbb{R}$. It is not hard to check from (39) that, for all $x\in[-c_{1},\infty)$, $|G_{\varepsilon}(x)|\leq c$ and $F_{\varepsilon}(x)\geq\log|x|-c-\log_{+}(c_{0}\varepsilon^{2}x)$. Hence
$B_{\varepsilon}(t)\leq c\int_{-c_{1}}^{\infty}\mathrm{d}x\,\frac{e^{-tx}}{(\log(|x|/c_{2})-\log_{+}(c_{0}\varepsilon^{2}x))_{+}{}^{2}}\ ,$ (43)
@@
$\int_{-c_{1}}^{c_{2}e^{3}}+\int_{c_{2}e^{3}}^{(c_{2}e^{3})\vee(1/t)}+\int_{(c_{2}e^{3})\vee(1/t)}^{1/c_{0}\varepsilon^{2}}+\int_{1/c_{0}\varepsilon^{2}}^{\infty}$ (44)
@@
$I^{\prime}_{2}=\frac{(x-c_{2}e^{3})}{(\log x/c_{2})^{2}}\Big{|}^{(c_{2}e^{3})\vee(1/t)}_{c_{2}e^{3}}+\int_{c_{2}e^{3}}^{(c_{2}e^{3})\vee(1/t)}\mathrm{d}x\,\frac{2\,(x-c_{2}e^{3})}{x(\log x/c_{2})^{3}}\ .$ (45)
On the right-hand side, bound the terms by $t^{-1}(\log((e^{3})\vee(1/c_{2}t)))^{-2}$ and $\frac{2}{\log e^{3}}I^{\prime}_{2}$, respectively. Doing so gives $I^{\prime}_{2}\leq c\,e^{t}\,t^{-1}|\log(t\wedge\frac{1}{2})|^{-2}$. As for $I_{3}$, since $x\leq 1/\varepsilon^{2}c_{0}$, the denominator of the integrand is $|\log(x/c_{2})|^{2}$. This is $\geq|\log(t\wedge\frac{1}{2})|^{2}/c$ because $x\geq(c_{2}e^{3})\vee(1/t)$. For $I_{4}$, the denominator is $(\log\varepsilon^{2}-c)^{2}$, which is $\geq|\log(t\wedge\frac{1}{2})|^{2}/c$ because $t\geq\varepsilon^{2}c_{0}$. Bounding the denominators this way and releasing the range of integration to $x\in[0,\infty)$ give $I_{3}+I_{4}\leq c\,t^{-1}|\log(t\wedge\frac{1}{2})|^{-2}$.